### v1

# Photon Number Coherence in Quantum Dot-Cavity Systems can be Enhanced by Phonons

Abstract
Semiconductor quantum dots are a versatile source of single photons with tunable properties to be used in quantum-cryptographic applications. A crucial figure of merit of the emitted photons is photon number coherence (PNC), which impacts the security of many quantum communication protocols. In the process of single-photon generation, the quantum dot as a solid-state object is subject to an interaction with phonons, which can therefore indirectly affect the PNC. In this paper, we elaborate on the origin of PNC in optically excited quantum dots and how it is affected by phonons. In contrast to the expectation that phonons always deteriorate coherence, PNC can be increased in a quantum dot-cavity system due to the electron-phonon interaction.
I Introduction
Quantum dots (QDs) have been shown to be excellent single-photon sources for many applications and have steadily been improved regarding certain figures of merit, such as purity,[1, 2, 3, 4, 5, 6] indistinguishability [6, 7, 8] and brightness.[2, 3, 4, 6] They have, in particular, been used for quantum-cryptographic applications,[9, 10] where the coherence between different photon-number states, known as photon number coherence (PNC) turns out to be a crucial figure of merit.[3] Some protocols, such as the famous BB84 key distribution protocol,[11] benefit from low coherence,[12, 13] while others benefit from the presence of PNC,[3, 14] such as twin-field quantum key distribution with single photons.[15] PNC has so far mostly been investigated using Poisson sources,[12, 13, 16] where usually a separate phase scrambling step is performed,[17, 18] which additionally decreases the secure key rate and can enable unambiguous state discrimination attacks, if it is performed imperfectly.[19, 13] Additionally, Poisson sources must be used at very low brightness to work in the few-photon regime,[20] while QDs provide photons on-demand [21] promising a significantly higher communication rate.[22]
A drawback of QDs as photon emitters is often their interaction with the solid-state environment, in particular with phonons. Phonons are known to induce dephasing in QDs, i.e, they typically destroy coherence,[23, 24, 25, 26, 27, 28, 29] lead to reduced excited state preparation [30, 31] and deteriorate photonic properties.[32] For PNC, the influence of phonons is less obvious, as phonons do not act on the photons directly, but only affect the QD coherence.
In this paper, we study the influence of phonons on the PNC in a laser driven QD-cavity system. After establishing the electronic coherence as the origin of PNC, we discuss losses and quantify the phonon influence up to temperatures of $\sim 100$ K. The typical expectation is that phonons will destroy coherence, as known from the electronic coherence. In contrast to this expectation, we find that for certain parameters, phonons can drastically change the expected behaviour of PNC and even increase it compared to the phonon-free case.
II The QD-cavity system
We consider an InGaAs/GaAs QD which we treat as a two-level system and is embedded in a mircocavity and is optically driven by a laser pulse. The uncoupled energies of the QD and the cavity are given by
$\hat{H}^{\text{QD}+\text{cav}}=\hbar\Delta\omega_{XL}|X\rangle\langle X|+\hbar\Delta\omega_{CL}\hat{a}^{\dagger}\hat{a},$ (1)
in the rotating frame of the laser. The state $|X\rangle$ describes the exciton and has the energy $\omega_{X}$. $|G\rangle$ denotes the QD ground state, the energy of which is set to zero, and the energy $\Delta\omega_{XL}=\omega_{X}-\omega_{L}$ is the energy difference between the exciton and the laser. Similarly, $\Delta\omega_{CL}=\omega_{C}-\omega_{L}$ describes the difference between cavity mode and laser energy. $\hat{a}$ ($\hat{a}^{\dagger}$) is the photonic annihilation (creation) operator of the cavity, which is coupled to the QD with the coupling strength $\hbar g=0.05~{}\text{meV}$,[33, 34, 35] by [36]
$\hat{H}^{\text{int}}=\hbar g\left(\hat{a}^{\dagger}|G\rangle\langle X|+\hat{a}|X\rangle\langle G|\right).$ (2)
We label the QD-cavity states $|S\,n\rangle$, where $n\in\mathbb{N}_{0}$ refers to a photonic state with $n$ photons and $S\in\{G,X\}$ signifies the QD state. The cavity-exciton detuning is $\Delta\omega_{CX}=\omega_{C}-\omega_{X}$. A classical laser pulse is used to excite the QD. It is described by the time-dependent pulse-envelope function $f(t)$ and adds the term [36]
$\hat{H}^{\text{laser}}=-\hbar\frac{f(t)}{2}\left(|G\rangle\langle X|+|X\rangle\langle G|\right)$ (3)
to the Hamiltonian. Gaussian pulses of the form
$f(t)=\frac{\Theta}{\sqrt{2\pi}\,\sigma}\text{e}^{-\frac{t^{2}}{2\sigma^{2}}},$ (4)
with a full width at half maximum FWHM=22ln $\text{FWHM}=2\sqrt{2\text{ln}\,2}\,\sigma=3~{}\text{ps}$ of the laser amplitude, and a pulse area $\Theta$, are used. We include the coupling to longitudinal acoustic (LA) phonons, that are known to be the most important ones for typical QDs.[30, 31] The LA phononic environment, where $\hat{b}_{j}$ ($\hat{b}^{\dagger}_{j}$) annihilates (creates) a phonon of mode $j$ with energy $\omega_{j}$, is described by the Hamiltonian [37, 38]
$\hat{H}^{\text{phon}}=\hbar\sum_{j}\omega_{j}\hat{b}^{\dagger}_{j}\hat{b}_{j}+\hbar\sum_{j}\left(\gamma_{j}^{X}\hat{b}_{j}^{\dagger}+\gamma_{j}^{X^{*}}\hat{b}_{j}\right)|X\rangle\langle X|,$ (5)
coupling each mode to the exciton with the coupling constant $\gamma_{j}^{X}$.[27, 28] Here, we consider a QD with a size of 3 nm. The phonon parameters of the InGaAs/GaAs QD were chosen as in Ref. [39]. Thus the total Hamiltonian is given by
$\hat{H}=\hat{H}^{\text{QD}+\text{cav}}+\hat{H}^{\text{int}}+\hat{H}^{\text{laser}}+\hat{H}^{\text{phon}}.$ (6)
Figure 1: Energy diagram of the most relevant states of the QD-cavity system $|G0\rangle$, $|G1\rangle$ and $|X0\rangle$. The latter two states are coupled via the cavity coupling $g$, while $|G0\rangle$, $|X0\rangle$ are coupled via the laser. Loss channels are indicated by blue arrows.
Radiative losses of the QD and cavity losses are described by Lindblad-superoperators [40, 41] affecting the density operator $\hat{\rho}$
$\mathcal{L}_{\hat{O},\delta}[\hat{\rho}]=\delta\left(\hat{O}\hat{\rho}\,\hat{O}^{\dagger}-\frac{1}{2}\left[\hat{\rho},\hat{O}^{\dagger}\hat{O}\right]_{+}\right),$ (7)
where $\hat{O}$ is an operator, $\delta$ is the decay rate and $[\,\cdot\,,\,\cdot\,]_{+}$ the anti-commutator. The rates used for the radiative and cavity losses are $\gamma=0.001~{}\text{ps}^{-1}$ and $\kappa=0.577~{}\text{ps}^{-1}$, respectively. The dynamics are given by the Liouville-von Neumann equation
$\frac{\text{d}}{\text{d}t}\hat{\rho}=-\frac{\text{i}}{\hbar}\left[\hat{H},\hat{\rho}\right]+\mathcal{L}_{|G\rangle\langle X|,\gamma}\left[\hat{\rho}\right]+\mathcal{L}_{\hat{a},\kappa}\left[\hat{\rho}\right],$ (8)
[…]
The QD density matrix is calculated by tracing out the cavity’s and phononic degrees of freedom
$\rho_{\mu\nu}\coloneqq\langle\mu|\text{Tr}_{\text{cav}}\,\text{Tr}_{\text{phon}}(\hat{\rho})|\nu\rangle$ $\mu,\nu\in\{G,X\},$ (9)
and similarly, the reduced photonic density matrix is obtained by
$\rho_{ij}\coloneqq\langle i|\text{Tr}_{\text{QD}}\,\text{Tr}_{\text{phon}}(\hat{\rho})|j\rangle$ $i,j\in\mathbb{N}_{0},$ (10)
where $|j\rangle$ denotes a state with $j$ photons in the cavity.
We focus on the experimentally relevant regime for single-photon emission, where the total excitation number is limited. The most relevant states are the energetically lowest three levels $|G0\rangle$, $|G1\rangle$ and $|X0\rangle$ which are displayed in Figure 1. In our numerical simulations, states with up to $n=2$ photons are accounted for. Because under our excitation conditions, occupations of cavity photon states with more than one photon are orders of magnitude smaller than the single-photon state, we henceforth focus our discussion on the photon number coherence between the vacuum and single-photon states $\rho_{01}$.
III The origin of photonic coherence
The laser excitation leads to coherence between the ground state and the excited state, which we will refer to as $\rho_{GX}$. Using the Heisenberg equation of motion of the photonic coherence $\rho_{01}$, one finds
$\frac{\text{d}}{\text{d}t}\rho_{01}(t)=\left(-\text{i}\Delta\omega_{CL}-\frac{\kappa}{2}\right)\rho_{01}(t)-\text{i}g\rho_{GX}(t).$ (11)
The equation demonstrates that $\rho_{01}$ represents a driven oscillator with a free oscillation frequency $\Delta\omega_{CL}$, that is damped via the rate $\kappa/2$. Most importantly, the coherence is driven by the QD coherence $\rho_{GX}$, hence the electronic coherence $\rho_{GX}$ is the source of photonic coherence $\rho_{01}$.
The solution of Equation (11) is formally given by
$\rho_{01}(t)=-\text{i}g\int_{-\infty}^{t}\rho_{GX}(t^{\prime})\,\text{e}^{\left(-\text{i}\Delta\omega_{CL}-\frac{\kappa}{2}\right)(t-t^{\prime})}\text{d}t^{\prime},$ (12)
linking the photon coherence to all past QD coherences by a memory function that constitutes a Lorentzian filter with a peak transmission at $\hbar\Delta\omega_{CL}\approx 0~{}\text{meV}$ and a band width determined by $\kappa/2$. Note that Equation (12) remains exact when phonons couple to the QD because the QD-phonon interaction commutes with all cavity operators. Yet, phonons can still affect the PNC via the QD coherences $\rho_{GX}$.
Figure 2: a) Short-time and b) long-time dynamics of the system excited by a $\pi$-pulse. Displayed is the QD population $\rho_{XX}$ and electronic coherence $|\rho_{GX}|$, single-photon state population $\rho_{11}$ and the photonic coherence $|\rho_{01}|$ for a QD-cavity detuning of $\hbar\Delta\omega_{CX}=1~{}\text{meV}$. These calculations were done without phonons, but the coherence behavior does not change qualitatively when phonons are included.
The time dependence of the absolute values of the coherences $|\rho_{01}|$ and $|\rho_{GX}|$ alongside the populations $\rho_{11}$ and $\rho_{XX}$ are shown in Figure 2 for an excitation with a $\pi$-pulse and a QD-cavity detuning of 1 meV. The pulse $f(t)$ is included for reference. The calculations were done without phonons. When a finite temperature and phonons are included, the qualitative behavior does not change. Three different dynamical regimes can be identified in the time evolution: The dynamics during the pulse, the short-time dynamics after the pulse and the long-time decay.
During the pulse, the system inverts, i.e., the population $\rho_{XX}$ goes to $\sim 1$, while the electronic coherence $|\rho_{GX}|$ reaches its maximum and then decreases again. During this maximum of the QD coherence, according to Equation (11) its emission into the cavity is large, such that $\rho_{01}$ increases quickly as well, producing the largest photonic coherence while the pulse is still active.
Just after the pulse, the photonic coherence drops off while displaying oscillations. These oscillations occur due to the detuning between QD and cavity, which is identical to $\Delta\omega_{CL}$ in Equation (11), because resonant pulses are used for the excitation. Because the stated $g$ and $\kappa$ values determine the system to be in the weak-coupling regime, these oscillations are only observed when adding a finite QD-cavity detuning $\hbar\Delta\omega_{CX}$. Additionally, a drop-off occurs on top of the oscillation.
After roughly 30 ps, the system reaches its long-time regime, where the decay becomes monoexponential with a single decay rate $\lambda_{+}$.
The dynamics of the QD and photonic coherences after the pulse can be described by a simplified approximate model, which is presented in Appendix A. From this model we extract the short-time and long-time decay by a linear combination of two exponential functions
$\rho_{01}(t)=A_{+}\,\text{e}^{\lambda_{+}t}+A_{-}\,\text{e}^{\lambda_{-}t}.$ (13)
with $\lambda_{-}$ standing for the initial faster decay and $\lambda_{+}$ for the long-time slower decay. In the phonon-free model, simple analytical expressions can be found for both decay rates and the amplitudes $A_{\pm}$. When phonons are included, Equation (13) still holds up to $\sim 70~{}\text{K}$, but $A_{\pm}$ and $\lambda_{\pm}$ must be determined numerically. The dependence of the decay rates $\lambda_{\pm}$ on the QD-cavity detuning and on temperature is discussed in Appendix B.
IV Time-integrated PNC
In most quantum-cryptographic studies, the details of the dynamics are not of interest.[3, 12, 14] Instead, a single figure of merit is desirable. A well defined quantity is the time integral over the absolute value of the reduced photonic density matrix
$\tilde{\rho}_{ij}:=\int_{-\infty}^{\infty}|\rho_{ij}|\,\text{d}t,$ (14)
with $i,j$ denoting the photon state. For $i\neq j$, this quantity is proportional to the PNC of the outside mode resonant with the QD-cavity system, once the photon has been fully emitted. Note that due to the integration, this measure of PNC does not distinguish between the two limiting cases, where the coherence is low at any given time, or is initially large and then decreases quickly.
We can use the fact that the integral in the calculation of the PNC from Equation (14) can be broken up into two components over the intervals $(-\infty,t_{0})$ and $(t_{0},\infty)$, where $t_{0}$ denotes the end of the laser pulse. The integral over the latter can then approximately be calculated by
$\tilde{\rho}_{01}^{(t_{0},\infty)}\approx F(\Delta\omega_{CX},g)\,|\rho_{GX}(t_{0})|,$ (15)
where $F$ is some filter function derived out of the filtering in Equation (12) in combination with Equation (13). The derivation of this formula and an explicit analytical expression for $F$ can be found in Appendix A. Importantly, the filter function $F$ can not depend on any properties of the pulse. Instead, any dependence on, for example the pulse area, lies exclusively in the initial electronic coherence $|\rho_{GX}(t_{0})|$. To understand the behavior of $\tilde{\rho}_{01}$, for which the integration in Equation (14) takes place over all times, it can be useful to consider only $\tilde{\rho}_{01}^{(t_{0},\infty)}$, which containes the integration over most relevant times. Now, we are able to investigate the influence that changes in parameters, like the QD-cavity coupling or the pulse area, have by determining their influence on the filter function $F$ or the $|\rho_{GX}(t_{0})|$. This simplifies the analysis greatly.
IV.1 Rabi rotations
Figure 3: Time-integrated absolute values of a) the one photon population and b) the coherence between 0 and 1 photons, displayed in dependence on the laser pulse area $\Theta$ for different temperatures $T$. The chosen temperatures are 4.2 K as well as 10 K to 110 K in 20 K steps from blue to red. The phonon-free case is displayed in green. All calculations were done with QD-cavity detuning $\hbar\Delta\omega_{CX}=0~{}\text{meV}$.
Figure 3 shows the integrated one-photon population $\tilde{\rho}_{11}$ along side the PNC $\tilde{\rho}_{01}$ as a function of pulse area $\Theta$ for different temperatures starting from $4.2$ K and then increasing from $10$ K in steps of $20$ K to $110$ K. For temperatures above 8080~{}8080 K optical phonons are expected to play a prominent role,[44] restricting the validity of our model to temperatures $T\lesssim 100$ K. The phonon-free case is displayed in green for reference.
In the two-level system, without a cavity or any rate decay, varying $\Theta$ produces Rabi rotations in the QD population, meaning the QD is fully excited (relaxed) when $\Theta$ is an odd (even) multiple of $\pi$. In turn, the coherence $\rho_{GX}$ is, with an absolute value of 0.5, at its largest for $\Theta=\left(n+\frac{1}{2}\right)\pi,n\in\mathbb{N}_{0}$, where the QD is half excited, and zero, whenever the QD is fully excited or fully relaxed. Considering this behavior of the electronic coherence and the proportionality between the electronic coherence and PNC from Equation (15), this makes $\Theta$ a very effective tuning knob for PNC. Some deviations exist, in large parts due to the influence of the cavity, but the phonon-free case and the low temperature cases in Figure 3 show the expected behavior. The deviations caused by the cavity are described in a model in Appendix C and produce for example the alternating height of the PNC maxima.
Phonons have several effects on the photon number and the PNC: A major effect is that the interaction with phonons leads to renormalization of the pulse area, shifting the laser intensity at which the QD is maximally excited towards larger values. Equally important is the effect that phonons dampen the Rabi rotations,[45, 46, 30, 23] resulting in a lower amplitude of the oscillations.
These effects can be observed in Figure 3. Focusing on the integrated occupation of the single photon state $\tilde{\rho}_{11}$ we can see both effects of renormalization and damping: The extrema shift towards higher pulse areas and the amplitude is decreased. For increasing temperature, the effective electron-phonon coupling becomes stronger and, correspondingly, the renormalization and damping becomes stronger.
These mechanisms similarly affect the electronic coherence and because of its close link to the photonic coherence, they also affect the PNC, which is shown in Figure 3  b). However, while the electronic coherence is largest for the $\pi/2$-pulse without phonons, it cannot increase when phonons are added. The PNC, in contrast can, since the filter function can have such an effect. As a result, it is possible for $\tilde{\rho}_{01}$ to increase for $\pi/2$ pulses above the phonon-free case. This surprising effect can be traced back to the renormalized coupling strengths,[29, 47, 34, 48] that influence this integrated quantity. Another remarkable effect is seen by looking at the behavior at pulse areas around $\pi$. While the PNC in the phonon-free case reaches almost zero, we find that the PNC remains much larger when phonons are accounted for. This effect is due to the damping of the Rabi rotations. Since a full inversion is not reached, the electronic coherence can still have finite values. At higher pulse areas, the PNC is then strongly damped and no more oscillations are visible, if the temperature is sufficiently large. Thus, depending on the pulse area, phonons can have coherence boosting or damping effects.
Figure 4: PNC for the pulse areas at which the QD occupation is maximized (red) and the PNC is maximized (orange) or minimized (blue), displayed for different temperatures. The respective phonon-free cases are displayed by the dashed lines. The calculations were done with $\hbar\Delta\omega_{CX}=0~{}\text{meV}$. The orange and blue lines are the maximal and minimal values from Figure 3 b). The unit of the PNC is ps, because it is the integrated value.
The impact of phonons on the first PNC extrema as a function a of temperature is summarized in Figure 4, where the maximal and minimal values of PNC are displayed in orange and blue, respectively. At low temperatures, both increase when phonons are included. As the temperature increases, the phonon-induced pure dephasing takes over and decreases PNC below the phonon-free value.
The most interesting curve is the red one, which is for the renormalized $\pi$-pulse, i.e., when the QD excitation is maximized just after the pulse. Interestingly, while the occupation has its maximum, the PNC is not at its minimum. Again, this is an effect of the cavity, which is explained by the model in Appendix C. When we therefore consider the PNC at the renormalized $\pi$-pulse, we first find an increase of the PNC up to about 60 K, and then the PNC decreases.
[…]
IV.2 QD-cavity detuning
Another important tuning parameter for the PNC is the QD-cavity detuning. Figure 5 a) displays the resulting time-integrated one-photon population $\tilde{\rho}_{11}$ and the 01-coherence $\tilde{\rho}_{01}$ for different QD-cavity detunings. $\tilde{\rho}_{01}$ increases by roughly one order of magnitude, when adding phonons and changes its shape. Whereas without phonons it has an almost imperceptible local minimum at $\hbar\Delta\omega_{CX}=0~{}\text{meV}$ and two maxima close to either side. The minimum is much more pronounced, when phonons are included.
To explain this phenomenon, we can again use the approximate model from Equation (15) and investigate the behavior of the filter function $F$ and the electronic coherence just after the pulse $|\rho_{GX}(t_{0})|$, separately. Figure 5 b) displays the filter function (orange), which possesses a central minimum for the QD-cavity coupling displayed. In the case without phonons, the electronic coherence just after the pulse (dashed green), has a central peak which is narrow compared to the minimum of $F$. As a result, the product of the two, which approximates the PNC over the course of the emission, only has a very slight minimum in the middle, producing the phonon-free $\tilde{\rho}_{01}$ in subfigure a). The QD coherence at time $t_{0}$ is generally much larger than in the phonon-free case [49, 50, 45, 30, 46] , which is again due to the damping described in Section IV.1. As a consequence, the product of it and $F$, still possesses the pronounced central minimum.
As can be seen in Figure 5 a), in the case with phonons, $\tilde{\rho}_{01}$ is generally smaller for negative QD-cavity detunings, compared to positive ones, which is explained by phonon-induced transitions between QD and cavity. For negative detunings, phonons provide an incoherent and fast relaxation mechanism from the exciton to the cavity, which decreases the QD population and the coherence quicker. But while this leads to an increase of $\tilde{\rho}_{11}$, $\tilde{\rho}_{01}$ decreases due to the incoherent nature of the phonon-induced transition.
Figure 6: a) Dependence of the filter function $F$ of the detuning $\Delta\omega_{CX}$ and the QD-cavity coupling $g$ b)/c) Temporal integration of the photonic 01-number coherence, $\tilde{\rho}_{01}$, displayed for different temperatures $T$ and QD-cavity detunings $\hbar\Delta\omega_{CX}$. The chosen temperatures are 4.2 K as well as 10 K to 110 K in 20 K steps. The phonon-free case is displayed in green. The pulse area is chosen to b) maximize the QD population and to c) minimize the QD polarization after the pulse.
The temperature-dependence of the PNC can also be investigated using the filter function $F$. To do this, we need to first investigate the QD-cavity coupling dependence of $F$, which is displayed in Figure 6 a). In the limit $\hbar g\to 0~{}\text{meV}$, the filter function converges towards zero, since the cavity is hardly ever excited when such low coupling is present. For small finite coupling strengths, only a single central maximum of $F(\Delta\omega_{CX})$ exists. It turns into a central minimum, flanked by two maxima, if the coupling strength is increased above a certain threshold. It is well known that phononic effects of increasing temperatures decrease the QD-cavity coupling strength.[29, 47, 34, 48] The effect of this can be seen in Figure 6 b) and c), which display the PNC for different temperatures and QD-cavity detunings for a) pulse areas that maximize QD occupation just after the pulse ended and b) pulse areas that minimize electronic coherence just after the pulse ended.
These two plots show that the PNC for different temperatures is governed by the two effects already mentioned: The general behavior follows the one observed in Figure 4. For example, the $\pi$-pulse produces an increase for small increasing temperatures, but then drops after roughly 55 K, whereas the pulse minimizing electronic coherence decreases the PNC monotonically for increasing temperatures, but also experiences phonon-induced damping as discussed in Section IV.1. The other effect that temperature introduces is the aforementioned renormalizaton of the QD-cavity coupling, adding the central minimum for small temperatures, which disappeared independently from the pulse area, at roughly 55 K.
V Conclusions
We have discussed the influence of phonons on the PNC of a single photon emitted from a QD. We started by establishing that the QD coherence is the source of PNC for a single photon source. Thus, all influences affecting electronic coherence, like losses or phonons, impact the PNC.
We have analyzed the PNC as function of pulse area, which for the QD occupation leads to the well-established Rabi rotations. In the phonon-free case, the control of the electronic system is passed on to the PNC. This changes significantly in the case with phonons. Although phonons lead to a damping of the electronic coherence, the PNC can increase compared to the phonon-free case.
Because of the renormalization of the QD-cavity coupling, different temperatures produce qualitatively different behavior of PNC for different QD-cavity detunings. Crucially, for low temperatures, the PNC increases for all values.
Our study shows that the influence of phonons on the PNC is highly non-trivial, as the simple assumption that the PNC exclusively decreases with phonons due to decoherence, does not hold true and other phononic effects must be taken into account. As PNC is crucial for many quantum cryptography applications, it is essential to always consider the influence of the solid-state-environment, which might be detrimental, but also could turn out as helpful.[3]
This implies that the strength of phonon coupling in applications benefiting from low PNC, such as standard/decoy BB84 QKD, unforgeable quantum tokens, quantum coin flipping and quantum bit commitment, should be carefully monitored to ensure practical security. On the other hand, phonons could provide better PNC in applications which require a phase reference such as twin-field quantum key distribution.[15]
Appendix A: Simple model of the relaxation
It is helpful to develop an understanding of the phonon-free case to determine the impact of phonons on the resulting coherences. Because we find the most important photonic coherence to be the 01-coherence and because the coherence between $|G0\rangle$ and $|X0\rangle$ is much larger than the one between $|G1\rangle$ and $|X1\rangle$, we can simplify the massive Hilbert space by looking only at the photonic coherence and the QD coherence, which are then mostly given by
$\rho_{01}$ $=\rho^{G0}_{G1}+\rho^{X0}_{X1}$ (16)
$\rho_{GX}$ $\approx\rho^{G0}_{X0},$ (17)
where the short notation
$\rho_{S_{2}n_{2}}^{S_{1}n_{1}}=\langle S_{1}n_{1}|\text{Tr}_{\text{phon}}(\hat{\rho})|S_{2}n_{2}\rangle$ (18)
was used. Through the Liouville-von-Neumann equation Equation (8) these two coherences are coupled to each other. For times when the laser is not active and if the phonon-free case is considered, these equations read
$\frac{\text{d}}{\text{d}t}\left[\begin{array}[]{c}\rho_{GX}\\ \rho_{01}\end{array}\right]=\left[\begin{array}[]{cc}-\text{i}\Delta\omega_{XL}-\frac{\gamma}{2}&-\text{i}g\\ -\text{i}g&-\text{i}\Delta\omega_{CL}-\frac{\kappa}{2}\end{array}\right]\left[\begin{array}[]{c}\rho_{GX}\\ \rho_{01}\end{array}\right].$ (19)
The behavior of the system can then be predicted by solving this ordinary differential equation (ODE) and depends on only two initial values. Let $t_{0}>0$ be the time at which the laser intensity after the maximum is sufficiently small to provide these initial values $\rho_{01}(t_{0})$ and $\rho_{GX}(t_{0})$, then the solution of Equation (19) can be written in the form given in Equation (13) in the main text. From the solution of this ODE, we obtain the squared absolute photonic coherence
$\begin{split}|\rho_{01}(t)|^{2}&=\frac{1}{|\text{det}\,T|^{2}}\Big{(}A(\Delta\omega_{CX})\,\text{e}^{2\text{Re}\lambda_{+}(t-t_{0})}\\ &+2\,\text{Re}\left[B(\Delta\omega_{CX})\,\text{e}^{(\lambda_{+}+\lambda_{-}^{*})(t-t_{0})}\right]\\ &+C(\Delta\omega_{CX})\,\text{e}^{2\text{Re}\lambda_{-}(t-t_{0})}\Big{)},\end{split}$ (20)
Figure 7: Comparison between full simulation and best fit of linear combination of two exponential functions, as in Equation 13 in the main text, for different temperatures. These calculations were done using an initial value problem with a fully excited QD.
with
$A(\Delta\omega_{CX})$ $=|\rho_{01}(t_{0})-\rho_{GX}(t_{0})T_{+}|^{2}|T_{-}|^{2}$ (21)
$\begin{split}B(\Delta\omega_{CX})&=\Big{(}-|\rho_{GX}(t_{0})|^{2}T_{-}^{*}T_{+}-|\rho_{01}(t_{0})|^{2}\\ &+\rho_{GX}(t_{0})\rho_{01}^{*}(t_{0})T_{+}\\ &+\rho_{GX}^{*}(t_{0})\rho_{01}(t_{0})T_{-}^{*}\Big{)}\,T_{-}T_{+}^{*}\end{split}$ (22)
$C(\Delta\omega_{CX})$ $=|\rho_{01}(t_{0})-\rho_{GX}(t_{0})T_{-}|^{2}|T_{+}|^{2}$ (23)
and
$Q$ $=\sqrt{\left(\text{i}\Delta\omega_{CX}+\frac{\kappa-\gamma}{2}\right)^{2}-4g^{2}}$ (24)
$\text{det}\,T$ $=T_{+}-T_{-}=-\frac{\text{i}}{g}Q$ (25)
$\lambda_{\pm}$ $=\frac{-\text{i}(\Delta\omega_{XL}+\Delta\omega_{CL})-\frac{\kappa+\gamma}{2}\pm Q}{2}$ (26)
$T_{\pm}$ $=\frac{\Delta\omega_{CX}}{2g}-\frac{\text{i}}{2g}\left(\frac{\kappa-\gamma}{2}\pm Q\right).$ (27)
Figure 7 shows that such a model is applicable even when phonons are active up to about 70 K. In that case, the decay rates must be fitted as they are affected by e.g. phonon renormalizations of the QD-cavity coupling, a polaron shift and a pure dephasing rate.
[…]
In case accurate values for $\tilde{\rho}_{01}$ are required, $\sqrt{|\rho_{01}(t)|^{2}}$ can be integrated numerically. However, since the integral is dominated by the slow exponential decay after the oscillations have ceased, we can approximate
$|\rho_{01}(t)|^{2}\hskip 5.0pt\stackrel{{\scriptstyle t-t_{0}>\!\!>2/\kappa}}{{\approx}}\hskip 5.0pt\frac{A(\Delta\omega_{CX})\,\text{e}^{2\text{Re}\lambda_{+}(t-t_{0})}}{|\text{det}\,T|^{2}}.$ (28)
Eventually, integration with Equation (21) and using $|\rho_{01}(t_{0})-\rho_{GX}(t_{0})T_{+}|\approx|\rho_{GX}(t_{0})T_{+}|$, which applies because $|T_{+}|\sim 10$ and because the initial electronic coherence is larger than the photonic one, results in an approximate formula for the integrated coherence after the excitation. To shorten notation, we introduce
$\tilde{\rho}_{01}^{\mathcal{G}}:=\int_{\mathcal{G}}|\rho_{01}(t)|\,\text{d}t$ (29)
for arbitrary integration intervals $\mathcal{G}$. Because of Equation (14) $\tilde{\rho}_{01}(t_{0})=\tilde{\rho}_{01}^{(-\infty,\infty)}$ applies. The stated approximations result in
$\tilde{\rho}_{01}^{(t_{0},\infty)}\approx\underbrace{-\frac{\frac{1}{\text{Re}\,\lambda_{+}}}{\left|\frac{1}{T_{-}}-\frac{1}{T_{+}}\right|}}_{=:F}|\rho_{GX}(t_{0})|$ (30)
for the integral over the relaxation period. In the main text, Equation (30) was reproduced and discussed as Equation (15). The full integral is thus approximated by
$\tilde{\rho}_{01}\approx F|\rho_{GX}(t_{0})|+\tilde{\rho}_{01}^{(-\infty,t_{0})}.$ (31)
This result can now be analyzed by investigating the three components $F$, $|\rho_{GX}(t_{0})|$ and $\tilde{\rho}_{01}^{(-\infty,t_{0})}$. The latter two are obtained from the full simulation because they are relatively complicated to calculate since they depend on the laser envelope $f$. The factor $F$, however, describes the impact of the relaxation after the excitation and is therefore independent of the details of the pulse and expressed by the simple analytical formulas given above. It is displayed in Figure 5 b) and 6 a).
The other components are displayed as a function of $\Delta\omega_{CX}$ in Figure 8. Now the existence of the central minimum, when phonons are active, and the lack of one without phonons (red dashed line), can be explained. The factor $F$ has a minimum at $\hbar\Delta\omega_{CX}=0~{}\text{meV}$ and two maxima on either side. Because of Equation (30) this minimum is inherited by $\tilde{\rho}_{01}^{(t_{0},\infty)}$, with and without phonons, as can be seen in Figure 8 b). However, because the initial value $|\rho_{GX}(t_{0})|$ has a higher peak around $\hbar\Delta\omega_{CX}=0~{}\text{meV}$ and is overall smaller for the case without phonons, in that case, the minimum is less pronounced.
Interestingly, the integrated coherence during the pulse $\tilde{\rho}_{01}^{(-\infty,t_{0})}$ appears to largely be independent of the presence of phonons. Due to the large initial coherence $|\rho_{GX}(t_{0})|$, the integrated coherence during the relaxation is about one order of magnitude larger when phonons are involved, than $\tilde{\rho}_{01}^{(-\infty,t_{0})}$. Because of this, the minimum can also be found in $\tilde{\rho}_{01}$. When phonons are not active, the already small minimum of $\tilde{\rho}_{01}^{(t_{0},\infty)}$ essentially disappears when adding $\tilde{\rho}_{01}^{(-\infty,t_{0})}$ to it, such that the central minimum of $\tilde{\rho}_{01}$ is almost imperceptible when phonons are not present.
The full simulation results are also displayed in Figure 8. Without phonons the simple analytical model describes $\tilde{\rho}_{01}$ very well, however, with phonons, it applies to a good extend only to positive detunings. For negative $\Delta\omega_{CX}$ a large discrepancy is found that can be explained by phonon-induced decoherence. Also, the maxima appear closer to $\hbar\Delta\omega_{CX}=0~{}\text{meV}$ and the minimum is higher for the full simulation. This is likely an effect of the renormalized QD-cavity coupling $g$, because $F$ displays exactly these traits for smaller $g$.
How the factor $F$ behaves for decreasing QD-cavity coupling is displayed in Figure 6 a). For $\hbar g\approx 0.01~{}\text{meV}$ the two maxima at either side of the minimum at $\hbar\Delta\omega_{CX}=0~{}\text{meV}$ merge into a single maximum. As a result, the disappearing central minimum for increasing temperatures in Figure 6 b) and c) from Sec. IV.2 can be explained by the effective decrease in QD-cavity coupling.
Appendix B: PNC long-time decay rates
Figure 9: Absolute values of the real components of the exponential decay rates $\lambda_{\pm}$ of the photonic coherence $\rho_{01}$ as a function of QD-cavity detuning $\hbar\Delta\omega_{CX}$ for temperatures 4.2 K and 10 K to 70 K in 20 K steps displayed in colors of the blue-red spectrum obtained by fitting Equation (13) to the numerical results of the real time behavior. The case without phonons is displayed in green. a) Displays the long-time decay rate and b) shows the short-term decay rate that governs the initial drop-off.
The absolute values of the real parts of the decay rates from Equation (13) $\lambda_{\pm}$ are displayed in their dependence of the QD-cavity detuning in Figure 9. Both are generally negative and therefore decay the photonic coherence. These values were obtained by fitting the model developed in Appendix A to the simulated data, as it was also done in Figure 7.
The long-time decay rate $\lambda_{+}$ in subfigure a) shows a central maximum for all temperatures (blue-red spectrum) and without phonons (green). Naturally this is due to the more efficient coupling between QD and cavity at these small detunings. What may be surprising, however, is that the height of this maximum decreases with increasing temperature. We accredit this to the phonon-induced renormalization of the QD-cavity coupling,[29, 47, 34, 48] which was also considered in Section IV.2. In fact, similar results to the high temperature decay rates can be obtained, when plugging a smaller QD-cavity coupling $g$ into Equation 26 (not shown). An unintuitive result of this is the fact, that the long-time decay rate of the coherence decreases despite the increased effects of phonons. The general shape of a single maximum near zero-detuning is kept at all temperatures, regardless of the presence of phonons, but when phonons are active at low temperatures the curve becomes asymmetric due to phonon emission processes assisting the creation of a cavity photon in case of negative $\Delta\omega_{CX}$. For higher temperatures, the opposite process, where a phonon is absorbed, can take place as well,[51] such that the curve becomes more symmetric with rising temperatures.
Figure 9 b) shows the short-time decay rates $\lambda_{-}$. In the phonon-free case, which is again displayed in green, only a small dip at zero detuning can be seen. Apart from this, the rate is independent from the detuning. The addition of phonons produces not just the asymmetry also observed in subfigure a), but interestingly also changes the general shape. For the highest displayed temperature of 70 K this is particularly obvious. The central minimum instead turns into a very pronounced maximum, which decreases to zero for large detunings. This change in shape does not change the earlier observation that $\Lambda_{+}$ is the more important rate when calculating $\tilde{\rho}^{(t_{0},\infty)}_{GX}$, since even at 70 K and even at a QD-cavity detuning of 1.5 meV, there is still about one order of magnitude difference between the two rates. Thus, the short-time decay is not significant in our simple model of the integration when calculating $\tilde{\rho}^{(t_{0},\infty)}_{GX}$.
Appendix C: PNC Rabi Rotations
When calculating Rabi rotations, a puzzling detail was found in Sec. IV.1, where we observed that every other peak of PNC in its pulse-area dependence is slightly higher than the next peak. This Appendix takes a look at a simple model, which reproduces this finding. Because the effect occurs in the phonon-free case as well, phonons can be excluded from this model. Additionally, all Lindblad-rates are discarded, because their effects are not needed to understand the observed behavior. As a result, the QD-cavity system can be described by a pure state. Under the assumption that the pulse is short enough to obtain a negligible two-photon component, all states higher than an occupation of $n=1$ may be discarded as well. To simplify the calculation, $\hbar\Delta\omega_{CX}=\hbar\Delta\omega_{LX}=0~{}\text{meV}$ are used. The resulting stripped down model is schematically displayed in Figure 10. For simplicity, the laser is described as a rectangular pulse starting at $t=0~{}\text{ps}$ and ending at some $t_{0}$. Thus, the coupling between the QD’s ground and excited states is given by
$f=\frac{\Theta}{t_{0}}.$ (32)
Figure 10: Energy diagram of the simplest model reproducing the $\pi/2$-alternations in the Rabi rotations of PNC.
Let the state be given by
$|\Psi(t)\rangle=c_{0}(t)|G0\rangle+c_{1}(t)|X0\rangle+c_{2}(t)|G1\rangle+c_{3}(t)|X1\rangle,$ (33)
then the Schrödinger equation for the coordinates $c_{i}$, results in
$\frac{\text{d}}{\text{d}t}\left(\begin{array}[]{c}c_{0}\\ c_{1}\\ c_{2}\\ c_{3}\end{array}\right)=-\text{i}\left(\begin{array}[]{cccc}0&-\frac{f}{2}&0&0\\ -\frac{f}{2}&0&g&0\\ 0&g&0&-\frac{f}{2}\\ 0&0&-\frac{f}{2}&0\end{array}\right)\cdot\left(\begin{array}[]{c}c_{0}\\ c_{1}\\ c_{2}\\ c_{3}\end{array}\right).$ (34)
Since it will be useful for later Taylor expansions, the Hamiltonian matrix is expanded with $\frac{2f}{2f}$, which results in
$\frac{\text{d}}{\text{d}t}\left(\begin{array}[]{c}c_{0}\\ c_{1}\\ c_{2}\\ c_{3}\end{array}\right)=\text{i}\,\frac{f}{2}\left(\begin{array}[]{cccc}0&1&0&0\\ 1&0&-2r&0\\ 0&-2r&0&1\\ 0&0&1&0\end{array}\right)\cdot\left(\begin{array}[]{c}c_{0}\\ c_{1}\\ c_{2}\\ c_{3}\end{array}\right),$ (35)
where $r\coloneqq g/f$ has been newly defined. Since the laser $f$ is much larger than the QD-cavity coupling $g$, this represents a system of two Rabi-oscillators, which are weakly coupled. If one solves it and varies the time $t_{0}$, a beating effect would be observed. But here, we want to investigate the pulse-area dependence. This problem can be solved analytically, which gives the diagonal matrix of eigenvalues $S$ and the transformation matrix into the eigenbasis $U$ to be
$S$ $=-\text{i}\,\frac{f}{2}\,\text{diag}(d_{0},d_{1},d_{2},d_{3})$ (36)
$U$ $=\left(\begin{array}[]{cccc}A/d_{0}&B/d_{1}&-C/d_{2}&-D/d_{3}\\ A&B&-C&-D\\ A&B&C&D\\ A(2r+d_{0})&B(2r+d_{1})&C(2r+d_{2})&D(2r+d_{3})\end{array}\right)$ (41)
$d_{0}$ $\coloneqq-r+\sqrt{r^{2}+1}$ $d_{1}\coloneqq-r-\sqrt{r^{2}+1}$ (42)
$d_{2}$ $\coloneqq r+\sqrt{r^{2}+1}$ $d_{3}\coloneqq r-\sqrt{r^{2}+1}$ (43)
where $A$, $B$, $C$ and $D$ are normalizing constants of the eigenvectors. Since $U$ is real and its column-vectors are normalized, its adjunct matrix is given by its transpose matrix, $U^{\dagger}=U^{T}$. The integration of the ordinary differential equation should take place for the duration of the pulse $t_{0}$, such that the resulting state after the excitation is given by
$\textbf{c}=U\,\text{e}^{S\,t_{0}}U^{\dagger}\textbf{c}(0),$ (44)
where the initial state is the ground state $|\Psi(0)\rangle=|G0\rangle$, such that $\textbf{c}(0)=(1,0,0,0)^{T}$. Since $d_{0}=-d_{3}$ and $d_{1}=-d_{2}$, as well as $A^{2}=D^{2}$ and $B^{2}=C^{2}$ apply, some of the exponential functions in Equation (44) merge into sines and cosines. In the arguments of those, $\Theta=ft_{0}$ is used, resulting in the exact solution
$\begin{split}\textbf{c}(\Theta)&=2A^{2}\left(\begin{array}[]{c}\frac{1}{d_{0}^{2}}\,\text{cos}\left(\frac{d_{0}}{2}\Theta\right)\\ \\ \frac{\text{i}}{d_{0}}\,\text{sin}\left(\frac{d_{0}}{2}\Theta\right)\\ \\ \frac{1}{d_{0}}\,\text{cos}\left(\frac{d_{0}}{2}\Theta\right)\\ \\ \text{i}\,\left(\frac{2r}{d_{0}}+1\right)\,\text{sin}\left(\frac{d_{0}}{2}\Theta\right)\end{array}\right)\\ &+2B^{2}\left(\begin{array}[]{c}\frac{1}{d_{1}^{2}}\,\text{cos}\left(\frac{d_{1}}{2}\Theta\right)\\ \\ \frac{\text{i}}{d_{1}}\,\text{sin}\left(\frac{d_{1}}{2}\Theta\right)\\ \\ \frac{1}{d_{1}}\,\text{cos}\left(\frac{d_{1}}{2}\Theta\right)\\ \\ \text{i}\,\left(\frac{2r}{d_{1}}+1\right)\,\text{sin}\left(\frac{d_{1}}{2}\Theta\right)\end{array}\right)\\ \end{split}$ (45)
Clearly, it is not very intuitive to read anything out of Equation (45). Instead, the fact that $r\ll 1$ can be used for some approximations
$2\frac{A^{2}}{d_{0}^{2}}$ $\approx 2\frac{B^{2}}{d_{1}^{2}}\approx\frac{1}{2}$ (46)
$2\frac{A^{2}}{d_{0}}$ $\approx\frac{1}{2}(1-r)$ (47)
$2\frac{B^{2}}{d_{1}}$ $\approx-\frac{1}{2}(1+r)$ (48)
$2A^{2}\left(\frac{2r}{d_{0}}+1\right)$ $\approx\frac{1}{2}(2r+1)$ (49)
$2B^{2}\left(\frac{2r}{d_{1}}+1\right)$ $\approx\frac{1}{2}(-2r+1).$ (50)
Similarly, the exponents can be expanded to give
$\frac{d_{0}}{2}\Theta\approx\frac{\Theta-gt_{0}}{2}$ $\frac{d_{1}}{2}\Theta\approx-\frac{\Theta+gt_{0}}{2}.$ (51)
When these approximations are plugged into Equation (45), one obtains
$\begin{split}\textbf{c}(\Theta)&=\frac{1}{2}\left(\begin{array}[]{c}\text{cos}\left(\frac{\Theta-gt_{0}}{2}\right)+\text{cos}\left(\frac{\Theta+gt_{0}}{2}\right)\\ \\ \text{i}\,\left[\text{sin}\left(\frac{\Theta-gt_{0}}{2}\right)+\text{sin}\left(\frac{\Theta+gt_{0}}{2}\right)\right]\\ \\ \text{cos}\left(\frac{\Theta-gt_{0}}{2}\right)-\text{cos}\left(\frac{\Theta+gt_{0}}{2}\right)\\ \\ \text{i}\,\left[\text{sin}\left(\frac{\Theta-gt_{0}}{2}\right)-\text{sin}\left(\frac{\Theta+gt_{0}}{2}\right)\right]\end{array}\right)\\ &+r\left(\begin{array}[]{c}0\\ \\ \text{i}\,\left[-\text{sin}\left(\frac{\Theta-gt_{0}}{2}\right)+\text{sin}\left(\frac{\Theta+gt_{0}}{2}\right)\right]\\ \\ -\text{cos}\left(\frac{\Theta-gt_{0}}{2}\right)-\text{cos}\left(\frac{\Theta+gt_{0}}{2}\right)\\ \\ 2\text{i}\,\left[\text{sin}\left(\frac{\Theta-gt_{0}}{2}\right)+\text{sin}\left(\frac{\Theta+gt_{0}}{2}\right)\right]\end{array}\right).\end{split}$ (52)
It is easy to see that this object is identical to the typical Rabi rotations of a two level system for $g=0$, where only the top two elements would oscillate – one with a cosine and one with sine – and the two states corresponding to the excited cavity would stay empty. In the following, this expression is simplified, but no further approximations are made. Using the identity
$\begin{split}a\,\text{sin}(x+\alpha)+b\,\text{sin}(x+\beta)=\\ \sqrt{a^{2}+b^{2}+2ab\,\text{cos}(\alpha-\beta)}\,\text{sin}(x+\delta),\end{split}$ (53)
with
$\delta=\text{atan2}(a\,\text{sin}\,\alpha+b\,\text{sin}\,\beta,a\,\text{cos}\,\alpha+b\,\text{cos}\,\beta),$ (54)
and using
$\tilde{A}\coloneqq\left|\text{cos}\left(\frac{gt_{0}}{2}\right)\right|$ $=\frac{\sqrt{2+2\,\text{cos}(gt_{0})}}{2}$ (55)
$\tilde{B}\coloneqq\left|\text{sin}\left(\frac{gt_{0}}{2}\right)\right|$ $=\frac{\sqrt{2-2\,\text{cos}(gt_{0})}}{2},$ (56)
the linear combinations in Equation (52) can be rewritten as
$\textbf{c}(\Theta)=\frac{1}{2}\left(\begin{array}[]{c}\tilde{A}\,\text{cos}\left(\frac{\Theta}{2}\right)\\ \\ \text{i}\tilde{A}\,\text{sin}\left(\frac{\Theta}{2}\right)\\ \\ \tilde{B}\,\text{sin}\left(\frac{\Theta}{2}\right)\\ \\ -\text{i}\tilde{B}\,\text{cos}\left(\frac{\Theta}{2}\right)\end{array}\right)+r\left(\begin{array}[]{c}0\\ \\ \text{i}\tilde{B}\,\text{cos}\left(\frac{\Theta}{2}\right)\\ \\ -\tilde{A}\,\text{cos}\left(\frac{\Theta}{2}\right)\\ \\ 2\text{i}\tilde{A}\,\text{sin}\left(\frac{\Theta}{2}\right)\end{array}\right).$ (57)
Similar to this step, another identity
$\begin{split}a\,\text{cos}\,\alpha+b\,\text{sin}\,\alpha=\\ \text{sgn}(a)\sqrt{a^{2}+b^{2}}\,\text{cos}\left(\alpha+\text{atan}\left(-\frac{b}{a}\right)\right),\end{split}$ (58)
can be used to combine these sines and cosines further and avoid writing the $c_{j}$ as linear combinations. The resulting amplitudes $X,Y,Z$ and phase angles $\delta_{1},\delta_{2}$ and $\delta_{3}$ are only weakly dependent on changes in $\Theta$ within any $[n\pi,(n+2)\pi]$ interval and may thus, for the purposes of this investigations be considered constant, since the main point of this Appendix is to investigate the PNC’s behavior within such an interval. The result is
$\textbf{c}(\Theta)=\frac{1}{2}\left(\begin{array}[]{c}\tilde{A}\,\text{cos}\left(\frac{\Theta}{2}\right)\\ \\ \text{i}X(\Theta)\,\text{sin}\left(\frac{\Theta}{2}+\delta_{1}(\Theta)\right)\\ \\ -Y(\Theta)\,\text{sin}\left(\frac{\Theta}{2}+\delta_{2}(\Theta)\right)\\ \\ -\text{i}Z(\Theta)\,\text{cos}\left(\frac{\Theta}{2}+\delta_{3}(\Theta)\right)\end{array}\right).$ (59)
Now the PNC just after the pulse can be written as
[…]
where $\sigma$ and
$W\coloneqq\frac{1}{2}\sqrt{\tilde{A}^{2}Y^{2}+X^{2}Z^{2}+2\tilde{A}XYZ\,\text{cos}(\delta_{1}-\delta_{2}+\delta_{3})}$ (61)
Figure 11: Results of Appendix C’s calculations. The QD-cavity coupling was chosen as $g=0.05~{}\text{meV}$ and the pulse-length is $t_{0}=10~{}\text{ps}$.
result out of Equation (53) again. Thus, the coherence can be written as a sine, which has a small phase angle and an offset. As one would expect, the frequency with which the coherence oscillates is double that of the population. When taking the absolute value, the offset produces the difference between every other PNC peak. These results are plotted in Figure 11, which displays the coordinates $c_{i}$, the absolute QD coherence, which was calculated calculated via $|\rho_{GX}|=|c_{0}^{*}c_{1}+c_{2}^{*}c_{3}|$, the parameters $W$, $\sigma$ and $O$ from Equation (60) and PNC, which was calculated using the absolute value of
$\rho_{01}=W\,\text{sin}(\Theta+\sigma)+O.$ (62)
The coordinates $c_{0}$ and $c_{1}$ behave very similarly to the usual driven two-level model. The states with an excited cavity are – with the chosen parameters $gt_{0}=0.7596$ – much less occupied. They oscillate with the same frequency, but they posses shifted phases. These phases converge $\delta_{2}\to\pi$ and $\delta_{3}\to-\pi/2$ for large $\Theta$. The last phase very quickly converges against $\delta_{1}\to 0$. These phase differences, especially $\delta_{3}$, which converges the slowest, create the offset $O$ in Equation (60) and thus result in the observed phenomenon. Figure 11 c) may suggest that the offset decreases, out of which one might conclude that the effect decreases. It does not, since the amplitude $W$ decreases roughly as quickly. This can be seen in the fact that Figure 11 b) and d) display the coherences for $\Theta\in[50\pi,55\pi]$, where $O$ and $W$ have largely converged.
Acknowledgements
P.C.A.H. is grateful for the advice and guidance he received from Dr. Michael Cossacchi and Dr. Tim Seidelmann.
D.E.R acknowledges financial support from the German Research Foundation DFG through project 428026575 (AEQuDot). We are grateful for fruitful discussions with Y. Karli, F. Kappe, V. Remesh and G. Weihs from University of Innsbruck.